# Chao-Han Huck Yang1, Ramani Duraiswami2, Dinesh Manocha2, Rafael Valle1, Bryan Catanzaro1

### v1

@@ Appendix H AF3 Training Details
In this section, we present the training settings of our models across all 5 stages, each with specific configurations. Details are in Table 12.
@@ I.1 Neural Audio Codec
We utilize a fully causal convolutional neural audio codec for efficient streaming audio decoding, following kim2024efficient ; siuzdakvocos .
Encoder. Input audio is first resampled to 44.1 kHz. It is then converted into Short-Time Fourier Transform (STFT) parameters using a hop size of 8 and a window size of 32. This STFT representation is processed by an initial 1x1 convolutional layer to produce 384-dimensional hidden embeddings. Following this, the signal undergoes three downsampling stages. Each stage consists of three causal 1D-ConvNeXt blocks liu2022convnet ; siuzdakvocos followed by a strided convolutional layer for downsampling. These strided convolutional layers use a stride and kernel size of 8. Each such layer doubles the hidden dimension, except for the final one, which produces a 512-dimensional output. The encoded output sequence is 4096 times shorter than the raw waveform, corresponding to approximately 10.8 frames per second.
Quantization. The encoded output is quantized into audio tokens using Residual Vector Quantization (RVQ) lee2022autoregressive ; kumarhigh . The number of RVQ levels is set to 72.
Decoder. The decoder mirrors the encoder’s architecture symmetrically, employing 1D transposed convolutional layers for upsampling and causal 1D-ConvNeXt blocks. The final convolutional layer reconstructs the STFT parameters, which are then transformed back into a raw audio waveform via an inverse STFT (iSTFT) similar to Vocos siuzdakvocos .
Training. The codec is trained using a combination of adversarial training and a mel-spectrogram reconstruction loss, following methodologies from DAC kumarhigh .
@@ I.2 Text-to-Speech (TTS) Module
Architecture and Operation. The TTS module’s transformer decoder processes a sequence formed by concatenating subword text tokens (from the main AF3 model) and previously generated audio tokens. The resulting hidden states from the transformer serve as conditional input to a multi-layer perceptron (MLP) block. This MLP block then iteratively predicts progressively higher levels of the RVQ tokens, a technique inspired by kim2024efficient . In practice, we employ 4 iteration steps during inference. A key aspect contributing to the system’s simplicity and low latency is that the model is designed to generate an audio token whenever a text token is emitted by the AF3 model, without requiring explicit alignments between text and speech.
Training and Configuration. During training, the transformer decoder utilizes teacher-forcing with ground-truth audio tokens. The MLP block is trained to estimate the parameters of a mixture-of-gaussians distribution where the number of mixtures is 1024. The objective is to maximize the log-likelihood of predicting the cumulative RVQ token embedding, following kim2024efficient . The decoder-only transformer has a configuration similar to DiT-XL peebles2023scalable ; lee2025dittotts . The MLP block consists of 3 layers, totaling 644 M parameters.
@@ I.3 Training Data and Processing
The models are trained on a comprehensive dataset comprising LibriTTS-R koizumi2023libritts , LibriHeavy kang2024libriheavy , the English portion of Multilingual LibriSpeech pratap2020mls , and proprietary datasets, totaling approximately 100,000 hours of audio. To support long-form speech generation, audio segments from the same speaker are randomly concatenated to create training samples with durations ranging from 1 to 120 seconds.

### v2

@@ Appendix H AF3 Training Details
In this section, we present the training settings of our models across all 5 stages, each with specific configurations. Details are in Table˜12.
@@ I.1 Neural Audio Codec
We utilize a fully causal convolutional neural audio codec for efficient streaming audio decoding, following [61, 103].
Encoder. Input audio is first resampled to 44.1 kHz. It is then converted into Short-Time Fourier Transform (STFT) parameters using a hop size of 8 and a window size of 32. This STFT representation is processed by an initial 1x1 convolutional layer to produce 384-dimensional hidden embeddings. Following this, the signal undergoes three downsampling stages. Each stage consists of three causal 1D-ConvNeXt blocks [79, 103] followed by a strided convolutional layer for downsampling. These strided convolutional layers use a stride and kernel size of 8. Each such layer doubles the hidden dimension, except for the final one, which produces a 512-dimensional output. The encoded output sequence is 4096 times shorter than the raw waveform, corresponding to approximately 10.8 frames per second.
Quantization. The encoded output is quantized into audio tokens using Residual Vector Quantization (RVQ) [69, 66]. The number of RVQ levels is set to 72.
Decoder. The decoder mirrors the encoder’s architecture symmetrically, employing 1D transposed convolutional layers for upsampling and causal 1D-ConvNeXt blocks. The final convolutional layer reconstructs the STFT parameters, which are then transformed back into a raw audio waveform via an inverse STFT (iSTFT) similar to Vocos [103].
Training. The codec is trained using a combination of adversarial training and a mel-spectrogram reconstruction loss, following methodologies from DAC [66].
@@ I.2 Text-to-Speech (TTS) Module
Architecture and Operation. The TTS module’s transformer decoder processes a sequence formed by concatenating subword text tokens (from the main AF3 model) and previously generated audio tokens. The resulting hidden states from the transformer serve as conditional input to a multi-layer perceptron (MLP) block. This MLP block then iteratively predicts progressively higher levels of the RVQ tokens, a technique inspired by [61]. In practice, we employ 4 iteration steps during inference. A key aspect contributing to the system’s simplicity and low latency is that the model is designed to generate an audio token whenever a text token is emitted by the AF3 model, without requiring explicit alignments between text and speech.
Training and Configuration. During training, the transformer decoder utilizes teacher-forcing with ground-truth audio tokens. The MLP block is trained to estimate the parameters of a mixture-of-gaussians distribution where the number of mixtures is 1024. The objective is to maximize the log-likelihood of predicting the cumulative RVQ token embedding, following [61]. The decoder-only transformer has a configuration similar to DiT-XL [93, 70]. The MLP block consists of 3 layers, totaling 644 M parameters.
@@ I.3 Training Data and Processing
The models are trained on a comprehensive dataset comprising LibriTTS-R [64], LibriHeavy [59], the English portion of Multilingual LibriSpeech [95], and proprietary datasets, totaling approximately 100,000 hours of audio. To support long-form speech generation, audio segments from the same speaker are randomly concatenated to create training samples with durations ranging from 1 to 120 seconds.
Appendix J Qualitative Examples
Figure 14: Demonstration of AF3’s capabilities on an audio captioning task. We prompt AF3 with an unseen audio clip—extracted from the Superman 2025 trailer (https://www.youtube.com/watch?v=2woCZg5QdVE)—captured in the wild. The model accurately identifies and describes background sounds, spoken content, speaker turns, and transcriptions, demonstrating strong audio understanding. Beyond this example, AF3 supports significantly more complex reasoning tasks. We invite readers to explore these capabilities via our public demo: https://huggingface.co/spaces/nvidia/audio-flamingo-3.